Find the length of the altitude of an equilateral triangle whose side measures 6 feet.: Find the length of the altitude of an equilateral triangle whose side measures 6 feet.
Answer by nerdybill(1040)

When you see the term "equilateral triangle" it should immediately register that all three sides are equal!
And, that the altitude (h) bisects one side in half.
Since x = 6 feet (given in problem)
Draw the diagram above...
Now, you can apply Pythagorean's theorem:
h^2 + 3^2 = 6^2
h^2 + 9 = 36
h^2 = 27
h = sqrt(27)
h = sqrt(3*3*3)
h = 3sqrt(3)